Grade 12th passMechanics

If a bullet of mass m penetrates a thickness h of a fixed plate of mass M..If the plate was free to move ,what will be the thickness penetrated??

Shubham Gupta
10 Years agoGrade 12th pass

1 Answer

sudhir soni
8 Years ago
Assuming retardation force F to same in both cases, we have ,Retardation a = F/mUsing v^2 = u^2 -2as, we have ,0=u^2 -2(F/m)hSo , F =mu^2/2h